OpenStudy (anonymous):

Given the function f(x) = 2(x + 8), find x if f(x) = 12 2 40 −2 14

OpenStudy (anonymous):

Substitute 12 for x ..

hero (hero):

You mean for f(x)

OpenStudy (anonymous):

f(x) = (12) + 8)

OpenStudy (anonymous):

Well don't do that....

OpenStudy (anonymous):

f(x) = 2(x+8) = 12

hero (hero):

12 = 2(x + 8)

OpenStudy (anonymous):

Woops, yeah go with hero's

hero (hero):

I'm interested in seeing @Tayloresc solve this

OpenStudy (anonymous):

I'm so confused by this

hero (hero):

If you have f(x) = 2(x + 8) and they tell you f(x) = 12, then replace f(x) with 12: If you do that you'll have 12 = 2(x+8) Can you solve that for x?

OpenStudy (anonymous):

no

hero (hero):

Do you know how to add, subtract, multiply and divide?

OpenStudy (anonymous):

yes

hero (hero):

Well, you can do this. Start with dividing both sides by 2. What do you get afterwards?

hero (hero):

You can do this: \[\frac{12}{2} = \frac{2(x+8)}{2}\]

hero (hero):

Which is the same as \[\frac{12}{2} = \frac{2}{2}(x+8)\]